A pharmacists needs to obtain a 70% alcohol solution. How many ounces of a 30% alcohol solution must be mixed with 40 ounces of an 80% alcohol solution to obtain a 70% alcohol solution?
Found 2 solutions by checkley71, Paul:
Answer by checkley71(8403)

(40+X).7=(40).8+.3X OR 28+.7X=32+.3X OR .4X=4 OR X=10
PROOF .7(40+10)=40*.8+.3*10 OR 35=32+3 OR 35=35

Answer by Paul(988)

Let the solution be x for 30%
Let y=40-x be the solution for 80%
Equation:
30x+80(40-x)=70(40)
30x-80x=70(40)-80(40)
-40x=-400
x=10
40-10=30
Hence, about 10 ounces is needed for 30% and 30 ounces is needed ofr 80%.
Paul.